OpenStudy (anonymous):

Calculate the area of the region bounded by the graphs of the given equations. y = 1/2x2, y = -x2 + 6

OpenStudy (anonymous):

that why i also aske people to double check that they typed the right question

OpenStudy (anonymous):

yes, sorry

OpenStudy (anonymous):

now the limits of intergration are from -2 tro 2

OpenStudy (anonymous):

so, again all you have to do is take the intergral from 0 to 2 of 6-5x^2/2 and double it

OpenStudy (anonymous):

now, could you please help me?

OpenStudy (anonymous):

well you need to integrate that

OpenStudy (anonymous):

is this \[y=\frac{1}{2}x^2\] adn \[y=-x^2+6\]

OpenStudy (anonymous):

i guess smooth man is typing up a long response, so i will wait for him

OpenStudy (anonymous):

-x2+6 = 1/2x2 6 = 3/2x2 4 = x2 x = +-2 integral from -2 to 2 of 6-x2-1/2x2 = int(-2 to 2) 6-3/2x2 Antiderivative is 6x - 3/6x3 evaluate from -2 to 2. Or from 0 to 2 and then double the result. 6(2) - 1/2(2)^3 - (6(-2) - 1/2(-2)^3) = 12 - 4 - (-12 + 4) = 24 - 8 = 16

OpenStudy (anonymous):

Yes, Satellite. Those are the functions.

OpenStudy (anonymous):

\[2\int _0^2-x^2+6-\frac{1}{2}x^2dx\] yes?

OpenStudy (anonymous):

Yeah. That would give it to you, Satellite.

OpenStudy (anonymous):

or somewhat more elegantly \[2\int_0^2-\frac{3}{2}x^2+6dx\]

OpenStudy (anonymous):

\[2\int_0^26dx =2\times 2\times 6=24\] by inspection

OpenStudy (anonymous):

Um what.

OpenStudy (anonymous):

Oooh. Nevermind. I see.

OpenStudy (anonymous):

and \[2\times \int_0^2 -\frac{3}{2}x^2dx=2\times -\frac{x^3}{2}|_0^2=-2^3=-8\]

OpenStudy (anonymous):

Yes. =) So the sum is 16.

OpenStudy (anonymous):

so i guess answer is 24-8=16 joui?

OpenStudy (anonymous):

ok got it!

OpenStudy (anonymous):

thank you. i have cal 2 exam today( it is online) . i have lots of questions . can i ask ?

OpenStudy (anonymous):

yes, i don't know why i add 24+ 8 then 32. i solved it more than 10 times but maybe careless.

OpenStudy (anonymous):

knock yourself out, but best to post them fresh. and if you are taking an on line exam and want to cheat, here is fine but doing them yourself here http://www.wolframalpha.com/ pretty much guarantees you get the right result, assuming you are careful with typing.

OpenStudy (anonymous):

no. i had 60 questions as home work. i have 20 of them to ask. when i finish with it. i will go and take it myself i want to learn. i do not want to cheat.

OpenStudy (anonymous):

com on just admit you wanna cheat lol

OpenStudy (anonymous):

I admire that so hard. If you legit want to learn and understand, you can expect a lot of help from me.

OpenStudy (anonymous):

can i ask more questions here?

OpenStudy (anonymous):

Yeah absolutely. The thread tends to lag as it gets longer though, so it's probably worth posting a new question.

OpenStudy (anonymous):

i have posted two questions but i did not get any feed back. what should i do., i am completely new and i don't know what to do .